The OfficeJet Printer can copy Maria's Dissertation in 8 Minutes. The LaserJet can copy the same document in 14 minutes. If the two printers work together,how long will they take to copy the dissertation.
Answer by stanbon(75887)

The OfficeJet Printer can copy Maria's Dissertation in 8 Minutes. The LaserJet can copy the same document in 14 minutes. If the two printers work together,how long will they take to copy the dissertation.
------------------------------------------------------
OJ Printer DATA:
time = 8 min/job ; rate = (1/8) job/min
------------------------------
LJ Printer DATA:
time = 14 min/job ; rate = (1/14) min/job
------------------------------
Together DATA:
time = x min/job ; rate = (1/x) min/job
=========================================
Equation:
rate + rate = together rate
1/8 + 1/14 = 1/x
14x + 8x = 8*14
22x = 8*14
x = 5.09 minutes
==========================
Cheers,
Stan H.
==========================